OpenStudy (anonymous):

what is the solution set of |4x-1|=3

OpenStudy (radar):

4x-1 = 3 and 4x-1 = -3 4x=1/3 x=1/12 4x=-2 x=-1/2

OpenStudy (radar):

Ans is x =1/12, x=-1/2